12123333333333333333:

Angela plays soccer and golf for a total of 125 minutes every day. She plays soccer for 45 minutes longer than she plays golf. Part A: Write a pair of linear equations to show the relationship between the number of minutes Angela plays soccer (x) and the number of minutes she plays golf (y) every day. (5 points) Part B: How much time does Angela spend playing golf every day? Show your work. (3 points) Part C: Is it possible for Angela to have spent 80 minutes playing soccer if she plays for a total of exactly 125 minutes and plays soccer for 45 minutes longer than she plays golf? Explain your reasoning. (2 points)

jhonyy9:

any idea ?

12123333333333333333:

NOPE.. :( I need this done by 12pm D: or I faillllll

jhonyy9:

from these given details result x +y =125 yes?

12123333333333333333:

but what about the 45 minutes longer?

jhonyy9:

so what mean this ?

12123333333333333333:

y + 45 = x can i put that too?

12123333333333333333:

it says a PAIR of equations anyways...

jhonyy9:

exactly

12123333333333333333:

ok...next?

12123333333333333333:

i dont get this next part...ik that if I subtract 45 from x then I will get y but idk the exact number

jhonyy9:

just substitute what you know inside the first equation

jhonyy9:

x+y=125 y+45=x

jhonyy9:

and so calcule the value of x and the value of y

12123333333333333333:

????

jhonyy9:

x+y=125 y+45=x so y+45+y=125 yes ? solve this for y

12123333333333333333:

80!

jhonyy9:

2y=80

jhonyy9:

y = ?

12123333333333333333:

but why is 2y, 80??

jhonyy9:

bc. x+y=125 and y+45 = x

12123333333333333333:

OHHHHHH i get it. 40

jhonyy9:

yes ure and now the value of x

jhonyy9:

sure

jhonyy9:

x+y=125 x+40=125 x=?

12123333333333333333:

so is the soccer 80?

jhonyy9:

no golf is the y

12123333333333333333:

ahhhh algebra stinks!

jhonyy9:

and y=40

jhonyy9:

how much you need adding to 40 to get 125 ?

12123333333333333333:

85

jhonyy9:

perfect

jhonyy9:

soccer 85 golf 40

jhonyy9:

40+45=85

12123333333333333333:

YESSSS. so it isnt possible for her to play soccer for 80 mins to get to the exact total

12123333333333333333:

Wow thank you soo much!!!

jhonyy9:

np with my pleasure

12123333333333333333:

Are you good at Biology?

jhonyy9:

not really

12123333333333333333:

Okay. Thanks!!!

jhonyy9:

have a nice day bye

12123333333333333333:

You too! Thanks